# Improving LLM Personas via Rationalization with Psychological Scaffolds

### v1

@@ Appendix A PB&J prompts
In Table 5 and Table 6, we provide the prompts that we use to generate scaffolded rationales and predictions from both GPT-4 and Mistral 7B.
System Message:
A person can be described as follows:
<demographic information>
The person has the following judgements: 1. <judgement_1>
2. <judgement_2>
3. <judgement_3>
User Message:
No Scaffold
Judgement: <judgement_i>
Experiences
Judgement: <judgement_i>
Big 5 Personality Traits
4. Agreeableness (compassion, respectfulness, trust in others)
5. Neuroticism (tendencies toward anxiety and depression)
Judgement: <judgement_i>
Schwartz Theory of Basic Human Values
Judgement: <judgement_i>
Primal World Beliefs
Judgement: <judgement_i>
System Message:
You are the following person:
<demographic information>
You have the following opinions: 1. <judgement_1>+<rationale_1>
2. <judgement_2>+<rationale_2>
3. <judgement_3>+<rationale_3>
User Message:
Based on your demographic and opinion information above, which answer would you select for the above question?
Question: <question>
Answer choices: <choice>
Table 6: Prompts used to predict answers, given a persona: We use a common system message that includes a user’s demographic information and all prior judgments, along with generated rationales. The user message then includes the exact question and answer choices.
Appendix B Dataset Details
B.1 OpinionQA
@@ B.2 MovieLens
B.3 Human Pilot Dataset
@@
Appendix C Model Details
Config GPT-4 Mistral 7B
model GPT-4-0613 Mistral 7B 0.2 Instruct
Number of parameters: Unknown Number of parameters: 7 billion
Rationale Generation
new_tokens 256 256
temperature 1 1
seed 6 6
GPU N/A, openai api call 3 A100
Inferring time 2 hours 1 hour
Answer Generation
new_tokens 10 275
temperature 0 0
seed 6 6
GPU N/A, openai api call 3 A100
Inferring time 2 hours 1 hour
Table 7: Model Configurations for Rationale and Answer Generation
Appendix D Additional Discussion
D.1 Does Point-of-View (PoV) matter while building LLM personas?
Approach Rationale PoV Answer PoV Accuracy
Demographics + Judgments - first 32.05
Demographics + Judgments - third 49.63
[…]
Appendix E Significance Tests
Table 9: Statistical significance of PB&J’s improvements for both GPT-4 and Mistral 7B: Results from one-tailed independent t-tests comparing PB&J variants to the best-performing baseline. We report both user-wise and question-wise significance.
Appendix F Analysis of LLM generated explanations
Approach
% of rationales
that have
atleast one
demographic
mentioned ($\uparrow$)
Demographic
Keyword
Density
(# of demographic
keywords /
rationale) ($\uparrow$)
Avg. diversity
of 2-grams
in rationales
per user ($\downarrow$)
Avg. diversity
of 2-grams
in rationales per
judgment ($\downarrow$)
[…]
Table 10: Analysis of LLM-generated explanations across psychological scaffolds, measuring demographic mentions, keyword density, and 2-gram diversity: Results suggest that strong persona alignment does not necessarily require high demographic reliance or linguistic variation. All analysis on rationales generated by GPT-4.
To better understand how different scaffolds influence the content and structure of LLM-generated rationales, we analyze explanations along three key dimensions: demographic reliance, lexical diversity, and structural repetition (Table 10). The percentage of explanations mentioning demographics and the density of demographic keywords capture how often rationales explicitly reference user identity. The average 2-gram diversity per user and per judgment reflects the linguistic variability of explanations, with lower values indicating greater diversity (i.e., less repetition). This score is also known as self-repetition Salkar et al. (2022); Shaib et al. (2024).
[…]
Appendix G More examples of PB&J-generated rationales
@@
Appendix H More on Psychological Scaffolds
A rich body of research has extended, debated, and validated each of the psychological scaffolds presented earlier. Among these, Primal World Beliefs stands out as the most linguistically motivated framework, which may explain its strong performance in our experiments. As described in Clifton et al. (2019), researchers analyzed historical texts and over 80K tweets using topic modeling and concept extraction to identify statements about how people perceive the world. These statements were then categorized through expert coding, consultation with social scientists, and discussions with religious focus groups, leading to the identification of 26 fundamental “primal world beliefs". These beliefs encapsulate deep-seated assumptions individuals hold about the world, such as whether the world is inherently safe or dangerous, simple or complex, and abundant or limited. Expanding on this work, Vu et al. (2022) introduced a Latent Beliefs Model, which leverages transformer-based embeddings and a modified GPT-2 model to automatically infer latent dimensions of human beliefs from social media text. This data-driven discovery of worldviews underscores the linguistic basis of primal beliefs and their connection to naturally occurring human explanations.
Another well-established framework in social psychology is Schwartz Theory of Basic Human Values Schwartz (1992); Schwartz et al. (2012). This theory posits that human values—fundamental guiding principles—are organized along universal motivational dimensions that drive behavior. Schwartz identifies ten broad value categories, such as self-direction (independence of thought and action), benevolence (concern for others’ welfare), and power (desire for dominance or control). These values are structured in a circular model, where adjacent values are more compatible, and opposing values (e.g., security vs. stimulation) tend to be in tension. A key aspect of Schwartz’s values is their cross-cultural validation; extensive empirical studies have shown that these value dimensions hold across diverse populations, making them a robust framework for modeling user judgments in an LLM setting. Unlike primal beliefs, which describe broad worldviews, Schwartz’s values provide a structured way to infer decision-making tendencies and moral considerations, making them particularly useful for understanding user preferences in ethical or societal questions.
Appendix I Wasserstein Distance Metrics

### v2

@@ Appendix B PB&J prompts
In Table 10 and Table 11, we provide the final prompts that we use to generate scaffolded rationales and predictions from both GPT-4 and Mistral 7B, that are used to generate results displayed in Table 3 and Table 4.
Appendix C Dataset Details
C.1 OpinionQA
@@ C.2 MovieLens
C.3 Human Pilot Dataset
@@
Appendix D More examples of PB&J-generated rationales
@@
Appendix E Significance Tests
Table 8: Statistical significance of PB&J’s improvements for both GPT-4 and Mistral 7B: Results from one-sided paired t-tests comparing PB&J variants to the best-performing baselines. We report both user-wise and question-wise significance.
Appendix F Combining Psychological Scaffolds
Approach Accuracy
[…]
Table 9: Combining Scaffolded Rationales: Given that scaffolds are key to improve LM persona, we ask to what extent can scaffolded rationales help? We experiment with four varying settings where scaffolded rationales are combined.
Given that psychological scaffold-based rationales help improve LM personas, we also investigated settings where we combined rationales from multiple scaffolds for a user. On the OpinionQA subset containing the Human Written rationales, we concatenate the top two (Primal World Beliefs and Schwartz Theory of Basic Human Values), bottom two (Experiences and Big 5 Personality Traits) and all rationales (ConcatAll) as context for the LM persona. This leads to subpar performance; we posit that adding rationales from all scaffolds is too noisy for the LM to be able to select reasonable justifications to support user judgments. In order to mitigate this, inspired by Self-Consistency Wang et al. (2023), we consolidate rationales from all scaffolds into a single rationale (CombineAllToOne). An additional LM is used for this post hoc processing, where the LM is provided the following instructions: “For a given judgment, you will be provided multiple rationales for why this person holds this judgment. Your job is to consolidate these rationales into one concise rationale. If the rationales are not consistent with each other or present diverging perspectives, you are allowed to pick a perspective, or also allowed to keep multiple perspectives for that judgment, based on what you think best reflects the person.” While this improves over the ConcatAll setting, this is still not enough signal for the LM persona, unlike single scaffolds like Primal World Beliefs.
System Message:
A person can be described as follows:
<demographic information>
The person has the following judgements: 1. <judgement_1>
2. <judgement_2>
3. <judgement_3>
User Message:
No Scaffold
Judgement: <judgement_i>
Experiences
Judgement: <judgement_i>
Big 5 Personality Traits
4. Agreeableness (compassion, respectfulness, trust in others)
5. Neuroticism (tendencies toward anxiety and depression)
Judgement: <judgement_i>
Schwartz Theory of Basic Human Values
Judgement: <judgement_i>
Primal World Beliefs
Judgement: <judgement_i>
System Message:
You are the following person:
<demographic information>
You have the following opinions: 1. <judgement_1>+<rationale_1>
2. <judgement_2>+<rationale_2>
3. <judgement_3>+<rationale_3>
User Message:
Based on your demographic and opinion information above, which answer would you select for the question shown below?
Question: <question>
Answer choices: <choice>
User Message with Chain of Thought:
Based on the above list of opinions and the demographic information, what would you choose for the question shown below? Provide your answer in the following format - "Reason: <reason>, Answer: <answer>". Only answer amongst the provided options, nothing else. Do not abstain from answering.
Question: <question>
Answer choices: <choice>
Table 11: Prompts used to predict answers, given a persona: We use a common system message that includes a user’s demographic information and all prior judgments, along with generated rationales. The user message then includes the exact question and answer choices, with or without rationales.
Appendix G Model Details
Config GPT-4 Mistral 7B
model GPT-4-0613 Mistral 7B 0.2 Instruct
Number of parameters: Unknown Number of parameters: 7 billion
Rationale Generation
new_tokens 256 256
temperature 1 1
seed 6 6
GPU N/A, openai api call 3 A100
Inferring time 2 hours 1 hour
Answer Generation
new_tokens 10 275
temperature 0 0
seed 6 6
GPU N/A, openai api call 3 A100
Inferring time 2 hours 1 hour
Table 12: Model Configurations for Rationale and Answer Generation
Appendix H Additional Discussion
H.1 Does Point-of-View (PoV) matter while building LM personas?
Approach Rationale PoV Answer PoV Accuracy
Demographics + Judgments - first 32.05
Demographics + Judgments - third 49.63
[…]
Appendix I Analysis of LM generated rationales
Approach
% of rationales
that have
atleast one
demographic
mentioned ($\uparrow$)
Demographic
Keyword
Density
(# of demographic
keywords /
rationale) ($\uparrow$)
Avg. diversity
of 2-grams
in rationales
per user ($\downarrow$)
Avg. diversity
of 2-grams
in rationales per
judgment ($\downarrow$)
[…]
Table 14: Analysis of LM-generated explanations across psychological scaffolds, measuring demographic mentions, keyword density, and 2-gram diversity: Results suggest that strong persona alignment does not necessarily require high demographic reliance or linguistic variation. All analysis on rationales generated by GPT-4.
Context analysis of LM-generated rationales
To better understand how different scaffolds influence the content and structure of LM-generated rationales, we analyze the rationales along three key dimensions: demographic reliance, lexical diversity, and structural repetition (Table 14). The percentage of rationales mentioning demographics and the density of demographic keywords capture how often rationales explicitly reference user identity. The average 2-gram diversity per user and per judgment reflects the linguistic variability of rationales, with lower values indicating greater diversity (i.e., less repetition). This score is also known as self-repetition Salkar et al. (2022); Shaib et al. (2024).
[…]
Comparing of LM-generated rationales with human-written.
While we emphasize that PB&J generates plausible yet synthetic rationales that may not be the actual reasoning user by the user, we acknowledge in Limitations that this is indeed a problem with any synthetic data used in an LM pipeline. However, we compare and contrast rationales generated by PB&J with Human Written ones on the OpinionQA Human Written subset. Table 15 presents examples that fall into two broad categories: cases where human and LM rationales reflect similar reasoning, and cases where they diverge. In the former, LM rationales often elaborate on the human rationale, mapping it onto structured beliefs such as the Safe or Enticing World. In the latter, human rationales sometimes cite specific personal anecdotes or cultural references that are absent in LM generations, which tend to rely on general psychological framing.
Despite this variability, PB&J still benefits from the structure provided by scaffolds like Primal World Beliefs, often yielding rationales that are plausible and useful for prediction, even when the reasoning path differs from that of the original user. This highlights both the promise and limitations of structured, synthetic rationalization for modeling human behavior.
Example Type: Similar Reasoning
Selected Judgment:
The growing number of legal immigrants working in the U.S. does not contribute too much to economic inequality in the country.
Primal World Beliefs rationale:
I firmly stand on the opinion that the growing number of legal immigrants in the U.S. does not contribute significantly to economic inequality. Grounded in the Safe World Belief, I perceive the world as one of cooperation where diverse people can coexist harmoniously. Immigrants often fill roles in our workforce that are frequently overlooked, thus contributing to the economy and overall stability of our country. From the Enticing World Belief perspective, our world is fascinating and enriched by the presence of individuals of different cultures and skillsets. Legal immigrants bring unique perspectives and cultural norms that can actually enhance our socioeconomic environment. Finally, looking through the lens of the Alive World Belief lens, I understand that people immigrate to the U.S. for a better life, not to contribute to economic disparity. They serve a purpose in the economy, thereby contributing to the betterment of society as a whole. They are not the ones who create or increase economic inequality. Instead, it’s systemic policies and structures that need addressing.
Human Written rationale:
With or without them, capitalism is designed to make economic inequality.
Example Type: Similar Reasoning
Selected Judgment:
All in all, I favor the use of animals in scientific research.
Primal World Beliefs rationale:
My belief is rooted in the Enticing World belief. I see the world as a place of progress and innovation. The use of animals in scientific research enables us to make strides in understanding diseases and developing treatments. While I acknowledge it’s a controversial topic due to ethical considerations, I believe the benefits outweigh the negatives. This kind of research promises great return, potentially leading to medical breakthroughs that can enhance human health and longevity.
Human Written rationale:
Although I do not endorse animal cruelty or abuse, I think some form of scientific research using animals is necessary for the greater human good. Even though these animals might suffer or die, they are helping to advance human society.
Example Type: Not Similar Reasoning
Selected Judgment:
Medical treatments today are considered worth the costs because they enable individuals to live longer and enjoy better quality lives.
Primal World Beliefs rationale:
I hold this opinion mainly because of my high score on the safe world belief. I believe in a world of cooperation and comfort, where people come together to ensure the wellbeing of everyone. Today’s medical treatments, despite their costs, allow individuals to live longer, healthier lives, which contributes to a safer, more comfortable world. I also have a high enticing world belief because medical advancements hold the possibility of new discoveries, further enhancing our quality of life and making the world more enticing. Lastly, my moderate alive world belief stresses the significance of purpose in every happening, and I see the purpose of these treatments in promoting longer and healthier lives.
Human Written rationale:
I loved listening to A Prairie Home Companion on NPR. A while back, the host, Garrison Keillor, had a heart valve replacement. Many of his ancestors had just died at his age, but he got to entertain us for another 20 and is still touring at 82.
Example Type: Not Similar Reasoning
Selected Judgment:
Medical treatments today often create as many problems as they solve.
Primal World Beliefs rationale:
I believe that we are living in a Safe World, where medical technologies are advancing rapidly. However, sometimes, this progression can be seen as a double-edged sword. While it offers the possibility to treat illnesses, it can also lead to unforeseen side effects and complications, potentially causing more harm than good. This is a result of the complexity of the human body and the inherent risks associated with medical treatments. We must balance the desire for healing with an understanding of the potential risks involved.
Human Written rationale:
Over priced medical care
Table 15: We provide examples of Primal World Beliefs and Human Written rationales for the same judgment and user. We observe that sometimes, LM-generated rationales reflect the same reasoning conveyed by the human-written ones, however, in cases where the human-written rationales are very niche, it is hard for the LM-generated rationales to mimic the same reasoning.
Appendix J Wasserstein Distance Metrics